OpenStudy (anonymous):

y = x - 2 7y - 4x = 10

OpenStudy (anonymous):

Do you have to find using elimination using multiplication?

OpenStudy (anonymous):

y = x - 2 7y - 4x = 10 Use addition or substitution to find the value of y for this set of equations.

OpenStudy (anonymous):

first equ can be re written as 7y=7x-14 or 7y-7x=-14 7y-7x=-14 7y-4x=10 (subtracting the above two equation) -3x=-24 x=8 now y=x-2 y=8-2=6 thus x=8 and y=6

OpenStudy (anonymous):

Use addition or substitution to find the value of y for this set of equations. x + y = 14 x = 6y

OpenStudy (texaschic101):

I would use the substitution method...You already know that y = x - 2. Now substitute x - 2 in for y in the 2nd equation. :)